OpenStudy (anonymous):

how to you solve 16^x - 5 * 8^x = 0 using logarithms?

OpenStudy (anonymous):

please show working out, thanx :)

OpenStudy (nikvist):

\[16^x-5\cdot 8^x=0\] \[2^{4x}-5\cdot 2^{3x}=0\] \[2^{3x}\cdot (2^{x}-5)=0\] \[2^{x}=5\] \[x=\log_{2}5\]